Example of Reductionism and Bracketing

Greg Hudson

September 30, 1994

I will examine the phenomenon of writing a computer program from both a reductionist and a phenomenological perspective. From the reductionist perspective, you may see a person typing on a keyboard connected to a computer with a disk and a monitor. The keyboard has keys with certain physical properties (e.g. they require a certain amount of force to depress, and return to position with a certain springiness). When keys are depressed, they cause certain electrical impulses to appear in the computer’s memory and they cause the pattern of pixels on the screen to change. While this description may help determine why programmers often get carpal tunnel syndrome, it does not tell you anything about computer programs, or even distinguish computer programming from writing a paper.

Using bracketing to describe this phenomenon, we first note a relationship between the general act of writing a computer program (a feature which is unthematically copresent) and the computer program in particular that is being written. Neither of these functionalities “are” anywhere (though there may be a representation of part of the program “in” the computer’s disk, or “in” the programmer’s mind); they depend on the context of the motivation for writing the computer program and other features. The functional structure of a computer program has many layers: for example, there is a relationship of Fundierung between the program and the syntactic parse tree of the program, between the parse tree of the program and the lexical tokens used to write the program, between the lexical tokens and their symbolic representations (letters and punctuation), between the symbolic representation of the tokens and their representation in the computer’s memory or on its disk. Note that below the level of the program, any of the factivities could be replaced with a different factivity; for instance, the syntax tree of the program could be represented as a diagram rather than as lexical tokens. Lastly, it is the motivated viewing of a program as a program that makes it into a program; if no one cares to view a sequence of symbols as a computer program, then it is just a sequence of symbols.